OpenStudy (anonymous):

Gravitational potential energy! A small confusion If i find out the work done by gravity in bringing a mass from infinity to a point R from M then that work should turn out to be positive but when i do the integral in the systematic way as follows i end up with negative answer :-/

OpenStudy (anonymous):

So here goes  Since the displacement is downwards and the force of gravity is ALSO downwards.. the work is positive and hence \[dw= +Fdx\] hence\[w = \int\limits_{\infty}^{R}{\frac{GMm}{x^2}dx}\] that yeilds \[w = -\frac{GmM}{R}\] so what is wrong in this?!?